Circular No. 7669

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

WZ SAGITTAE R. Ishioka, M. Uemura, K. Matsumoto, and T. Kato, Kyoto Univer-

sity; K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka, Kyushu University, on behalf of the VSNET collaboration team and the Variable Star Observers League in Japan, report: “WZ Sge is undergoing its first outburst since 1978 (cf. IAUC 3311), as indicated by the following visual magnitude estimates: July 23.565 UT, 9.7 (T. Ohshima, Gifu, Japan); 23.597, 9.4 (H. Itoh, Tokyo, Japan); 23.600, 9.6 (H. Maehara, Saitama, Japan); 23.622, 9.8 (T. Watanabe, Shizuoka, Japan); 23.716, 8.6 (Itoh); 23.843, 8.7. CCD time-resolved photometry at Kyoto University on July 23.6–23.8 showed a rapid brightening from mag 9.7 to 8.7 within 5 hr. Short-term modulations with timescales of 0.01–0.03 day were superimposed on the rapidly rising trend. These modulations will grow into ‘early superhumps’, which are observed in early stages of the outburst of WZ Sge-type stars (cf. IAUC 6164, 6517, 7552). A preliminary spectrum (range 450–670 nm) of WZ Sge on July 23.74 (taken by Ayani with the BAO 1.01-m telescope) shows Hα and Hβ lines in absorption on a blue continuum. This feature differs from one at 1978 outburst (Brosch et al. 1980, Ap.J. 236, L29; Oltorani et al. 1980, A.Ap. 87, 31), when the Hα line was observed as emission (on the day that the outburst was detected) and became stronger in its intensity with time.”

J. A. Mattei, AAVSO, also reports that, according to observations in the AAVSO International Database, the last outburst of WZ Sge occurred on 1978 Dec. 1, when the star reached mag 8.0; it slowly faded to mag 10.8 by Dec. 30, with prominent superhump oscillations of several tenths of a magnitude throughout. On 1978 Dec. 31, the variable faded abruptly to mag 12.5 and then recovered to 11.1 by 1979 Jan. 3, before slowly fading to minimum by Apr. 1979. Earlier outbursts of WZ Sge include 1913 (when it reached m_{pg} = 7.0) and 1946 (when it reached m_{v} = 8.0). Recent visual magnitude estimates of WZ Sge submitted to the AAVSO and/or to the Central Bureau: July 20.061 UT, 15.4 (G. Poyner, Birmingham, England); 22.969, 15.3 (M. Reszelski, Szamotuly, Poland); 23.030, [13.6 (H. McGee, West Clandon, Surrey, England); 23.876, 8.1 (C. Jones, Laindon, Essex, England); 23.886, 8.4 (P. Schmeer, Bischmisheim, Germany); 23.895, 8.6 (R. J. Bouma, Groningen, The Netherlands); 23.922, 8.6 (Poyner); 24.013, 7.9 (F. Vohla, Altenburg, Germany); 24.050, 8.1 (G. Comello, Groningen, Netherlands; 24.105, 8.0 (J. Bortle, Stormville, NY); 24.244, 8.2 (R. Royer, Lakewood, CA). CCD V magnitudes by J. D. West, Mulvane, KS: July 24.110, 7.95; 24.126, 8.41 (comparison star HD 191083, V = 8.76).

2001 July 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green